18 Un-Natural Laws

• Murphy’s Law: If anything can go wrong, it will.

• O’Tool’s Commentary on Murphy’s Law: Murphy was an optimist.

• Non-reciprocal Laws of Expectations: Negative expectations yield negative results. Positive expectations yield negative results.

• Howe’s Law: Every man/woman has a scheme that will not work.

• Zymurgy’s First Law of Evolving Systems Dynamics: Once you open a can of worms, the only way to re-can them is to use a larger can.

• Etorre’s Observation: The other line moves faster.

• Skinner’s Constant (Flannagan’s Finagling Factor): That quantity which , when multiplied by, divided by, added to, or subtracted from the answer you get, gives you the answer you should have gotten.

• Law of Selective Gravity: An object will fall so as to do the most damage.

• Jenning’s Corollary: The chance of the bread falling with the buttered side down is directly proportional to the cost of the carpet.

• Gordon’s First Law: If a research project is not worth doing, it’s not worth doing well.

• Maier’s Law: If the facts do not conform to the theory, they must be disposed of.

• Hoare’s Law of Large Problems: Inside every large problem is a small problem struggling to get out.

• Boren’s First Law: When in doubt, mumble.

• The Golden Rule of Arts and Sciences: Whoever has the gold makes the rules.

• Barth’s Distinction: There are two types of people: those who divide people into two types, and those who don’t.

• Ninety-Ninety Rule of Project Schedules: The first 90% of the task takes 90% of the time, and the last 10% takes the other 90%.

• Farber’s Fourth Law: Necessity is the mother of strange bed-fellows.

• Segal’s Law: A man/woman with one watch knows what time it is. A man/woman with two watches is never sure.